the product of two consecutive odd numbers is 195.find the sum of the two numbers
Answer by mananth(16949)

let the numbers be x & x+2
x(x+2)=195
x^2+2x-195=0
x^2+15x-13x-195=0
x(x+15)-13(x+15)=0
(x+15)(x-13)=0
x=-15 OR 13
x,y = -15,-13
x+y=-28
OR
x+y = 13+15 =28